OpenStudy (destinykiara99):

Find the inverse of the function. f(x) = x3 - 8 Not a one-to-one function f-1(x) = cube root of x plus eight f-1(x) = cube root of quantity x minus eight f-1(x) = cube root of quantity x plus eight

TheSmartOne (thesmartone):

To find the inverse of a function: change f(x) to y switch y to x and the x to a y solve for y

jhonyy9 (jhonyy9):

so y = x^3 -2^3

TheSmartOne (thesmartone):

It's x^3 when people copy paste the question \(x^3\) becomes x3

OpenStudy (cosmo):

\[y=x^3−8\] \[x=y^3−8\] \[x+8=y^3\] \[y=\sqrt[3]{x+8}\]